### v2

@@ 1 Definition
In general, robustness is a predicate that applies to a single entity. For instance, we might consider a sensor robust if it is resilient to disturbances from the environment. In more detail, robustness refers to the ability of a system, model, or entity to maintain stable and reliable performance across a broad spectrum of conditions, variations, or challenges, demonstrating resilience and adaptability in the face of uncertainties or unexpected changes. Hence, we use the following general definition for Machine Learning (ML) Model robustness.
ML Model robustness denotes the capacity of a model to sustain stable predictive performance in the face of variations and changes in the input data. To better understand this definition, we add concrete illustrations of how performance degradation and data changes manifest in real-world scenarios.
  • Examples of variations and changes in the input data:
      Variations in input features or object recognition patterns that challenge the inductive bias learned by the model from the training data.
      Production data distribution shifts due to naturally occurring distortions, such as lighting conditions or other environmental factors.
      Malicious input alterations that are deliberately introduced by an attacker to fool the model or even steer its prediction in a desired direction.
      Gradual data drift resulting from external factors, such as evolution in social behavior and economic conditions.
  • Examples of model flaws and threats to stable predictive performance :
      Exploitation of irrelevant patterns and spurious correlations that will not hold up in production settings.
      Difficulty in adapting to edge-case scenarios that are often underrepresented by training samples.
      Susceptibility to adversarial attacks and data poisonings that target the vulnerabilities of overparametrized modern ML models.
      Inability of the model to generalize well to gradually-drifted data, leading to concept drift as its learned concepts become obsolete or less representative of the current data distribution.
Nevertheless, leaving the range of input data changes unspecified makes it hard to assess the robustness of ML model in practice. We should define the data changes against which the model would be tested. Even if the naturally occurring data distribution shifts are often unanticipated, we usually come up with data distortions that can serve as a proxy for unforeseen data shifts and help us compare the robustness of different ML models. Furthermore, the objective of sustaining stable predictive performance is vague. It is often sufficient for the model to maintain its performance to a certain degree (i.e., the tolerance level) against unexpected changes in input data. This level of tolerance depends on the application context and the assurances needed. For example, the target tolerance for an ML model designed to support clinical decision-making is considerably lower than for an ML model that has been designed to detect spam emails.
@@
Robustness is a standalone epistemic concept, presupposing the generalizability of the model’s inductive bias on the in-distribution data, and extending further to evaluate the stability and resilience of its inductive bias in real-world deployment scenarios.
1 Robustness complements (iid) generalizability
In supervised learning, models are commonly estimated via empirical risk minimization (ERM) goodfellow2016deep , a principle that considers minimizing the average loss on observed samples of data, as an empirical estimate of the true risk, i.e., the expected true loss for the entire input distribution. ERM assumes that training and test data are identically and independently distributed (a.k.a. i.i.d. assumption), known as closed-world assumption. The i.i.d generalization refers to the ability of a trained model to deal with novel data inputs, but drawn from the same or close distribution as the training set, called in-distribution data (ID). The i.i.d. generalization ensures stable predictive performance under static environmental conditions, but it provides no guidance on how to handle out-of-distribution data (OOD) yang2021generalized . In contrast, robustness focuses on capturing the level of predictive performance maintained by the trained model in dynamic environment settings, where input data constantly changes. Robustness could be of little concern if the model fails to i.i.d. generalize well. In order to achieve robustness, we consider i.i.d. generalization to be a necessary but not sufficient condition. For instance, ML models might fail to i.i.d. generalize due to unreliable inductive bias (shortcut learning) or under-fitting (trained with too little/biased data). These models will most likely also perform poorly when input data distributions change.
The ascent of deep learning has initiated a new era in artificial intelligence (AI), empowering models to tackle open-world learning challenges in domains like face recognition guo2016ms and autonomous driving bojarski2016end . The term “generalization” has expanded beyond denoting the model’s performance strictly in an i.i.d. setting, encompassing its ability to cope with out-of-distribution situations, hereafter referred to as o.o.d. generalization. Although o.o.d. generalization seems to fit our definition of robustness, it actually refers to the overall predictive performance of an ML model beyond its in-data distribution. It lacks a proper definition of its scope and success criteria since there is no indication of which data distributions the model should generalize, or how strict the original predictive performance should be maintained. In contrast, ML model robustness is an inherently causal concept since it concerns two causally related entities: the level of predictive performance and the input data change domain, so its assessment, as a model property, requires detailed specification of both.
@@ 2 Robustness is a requirement of Trustworthy AI
The adoption of machine learning, especially deep neural networks, has been largely promoted as a result of its impressive performance in terms of accuracy. Meanwhile, a variety of challenges outside of accuracy expectations have emerged, such as malicious attacks against ML-powered systems and misuses of ML that could be harmful. As a result, the Artificial Intelligence (AI) trustworthiness standards liu2022trustworthy ; li2023trustworthy have been established to outline representative requirements for current AI systems. These encompass six critical facets: (i) Safety & Robustness, (ii) Nondiscrimination & Fairness, (iii) Transparency & Explainability, (iv) Privacy, (v) Accountability & Auditability, and (vi) Environmental Well-being. The complex interplay between these aspects is vital in fostering trustworthy real-world AI systems. For instance, maintaining data privacy might interfere with the desire to explain the system output in detail.
@@
In this sense, robustness is an integral part of AI trustworthiness, while interacting and combining with the other aspects. In fact, model robustness is a cornerstone of safety because robust AI systems are able to deal with unexpected inputs and perturbations without compromising their functionality. It is especially crucial to ensure the appropriate level of model robustness in safety-critical applications, where erroneous behaviors and failures can have catastrophic consequences.
Furthermore, the safety of AI systems relies on the integration of two pivotal elements along with robustness assurance: reliable quantification of uncertainty and effective out-of-distribution detection capabilities.
@@
The uncertainty quantification abdar2021review refers to methodologies for evaluating uncertainties associated with predictions made by a ML model. It involves assessing the confidence levels or lack thereof in the model’s predictions, taking into account factors such as data variance and model error. The quantified uncertainties combined with predictions enable more informed decision-making with AI. More specifically, uncertainty in predictions can manifest as either aleatoric (non-reducible), stemming from inherent randomness in the data, such as noise, or as epistemic (reducible), arising from limitations within the model’s bias and the training data used. In simpler terms, the uncertainty component provides the AI system with a way to “know what they do not know”. Its main contribution to safety is that uncertain predictions can be ignored from the decision-making flow, avoiding risks in real-world applications. Estimating uncertainties significantly benefits the model robustness analysis by providing an essential means to gauge how individual predictions respond to changes in their data points. This aids in delimiting the domain of input changes on which an ML model is expected to be robust.
@@
Out-of-distribution detection yang2021generalized involves identifying OOD instances at test time that differ significantly from the in-data distribution and might result in mispredictions. It also serves the same purpose of recognizing the boundaries within which the model’s patterns are applicable, avoiding the use of its predictions when such restrictions are violated. OOD detection strengthens the reliability of uncertainty quantification by filtering out unusual inputs, on which the uncertainties, as any statistically-inferred estimates, are unlikely to be dependable. Therefore, out-of-distribution detection also contributes to further refinement of the data changes domain during the model robustness analysis, and its deployment in safety-critical systems ensures effective operation of robust ML models.
@@ 3 Adversarial vs Non-adversarial robustness
Non-adversarial robustness studies the model’s ability to maintain its performance across data distribution shifts arising from naturally-occurring distortions or synthetic data variations that represent conditions more likely to occur in the real world. For instance, a natural shift in images of traffic signs, collected in an area where it rarely snows, can be images of the same sign under severe snowing conditions gojic2023non . Alternatively, a partial discoloration of the traffic sign image, i.e., a region replaced by white pixels, may mimic the effect of snow or other neutral obstructions. Natural data shifts often result from changing environmental conditions that lead to mismatch between the deployment and the training distribution xie2019multi . The changes can be both temporal (i.e., social norms, evolution of subjects’ behavior) or non-temporal (i.e., changes across locations, sensor settings). The non-adversarial robustness of the ML model ensures reliability across various real-world scenarios, including natural noise, or changes in inputs that might happen organically without malicious intent.
@@ 4 Robustness Measurement
In the following, we introduce several robustness scoring metrics developed by researchers to evaluate the stability of models’ performance when data changes drenkow2021systematic . These metrics are designed with the assumption of having two datasets: one clean, sampled from in-distribution data, and another perturbed, sampled from shifted or even out of distribution data.
@@
The more the robustness score of a model is close to one, the more it is robust to the considered perturbation. To measure the robustness score of a neural network, $m$, to a set of perturbations $S$, we can use:
$R_{m}^{S}=\sum_{\phi\in S}R_{N}^{\phi}$
Furthermore, Hendrycks and Dietterich hendrycks2019benchmarking define robustness as average-case performance over a set of corruptions that gives rise to the definition of mean Corruption Error (mCE) and relative Corruption Error (rCE). The metrics mCE and rCE can be defined as:
@@
where $E$ is the classifier’s error rate, $\mathcal{C}$ is a set of corruptions, and $s$ is the severity of the corruption. The AlexNet model krizhevsky2012imagenet served as a common reference point among models.
Recently, Taori et al. taori2020measuring present two metrics for effective and relative robustness which considers how performance on natural distribution shifts relates to performance on an original test set. The metrics are defined as follows:
$aligned\rho(f)&=acc_{2}(f)-\beta\cdot acc_{1}(f)\\ \tau\left(f^{\prime}\right)&=acc_{2}\left(f^{\prime}\right)-acc_{2}(f)$
where $f$ is the model under test, $acc_{1}$, $acc_{2}$ are the accuracy on the original and shifted datasets respectively, $\beta$ is a log-linear fit to the baseline accuracy of a large set of independent models on the original (clean) test set, and $f^{\prime}$ represents the model resulting from a robustness improvement. The effective robustness measure captures how well a specific model does beyond what is expected given a group of models in general. The relative robustness measure quantifies the effect of a robustness amelioration strategy on the accuracy under distribution shift. Overall, a robust model should obtain both positive effective and relative robustness scores.
@@ 5 Robustness Indicators: Reproducibility and Explainability
We introduced robustness as a quality predicate for an ML model that should be carefully specified, evaluated, and sometimes certified before the deployment of the model in production. Nevertheless, the ML model is the result of a ML engineering process, thus some desirable properties in regards to this process should be satisfied in order to prepare the grounds to reach the level of robustness.
@@
A reproducible ML workflow controls its randomness and is resilient to slight variations in data samples, i.e, there are high chances to converge to the same conclusions when run on two samples from the same underlying distribution albertoni2023reproducibility . For researchers in all scientific fields, including machine learning, reproducibility is essential in order to achieve the same results with the same data and algorithms. The absence of reproducibility may lead scientists to claim gains from changing one parameter while the real source of improvement is a hidden source of randomness. Furthermore, reproducibility is also a strong stability property since it ensures a high probability of replicating results when the datasets are drawn from the same distribution. Consequently, achieving such property on a given learning problem ensures less bias in model estimation (M) since achieving such property confirms that the difference between the empirical risk of M obtained on the training samples and the (true) risk of M is marginal, which implies that M is more likely independent of the training set. For instance, LABEL:underspecificationMLmodel constructs an ensemble of predictors from a given model by perturbing components of the ML pipeline like random seeds for initialization, and retraining the model several times. As a result, they demonstrate that output differences between predictors against stress tests can be used as a conservative indicator of the ML pipeline credibility, i.e., its ability to produce a robust model.
@@
Explainability is crucial in the ML workflow, which addresses how an AI model makes decisions carvalho2019machine . Being aware of the reasoning behind predictions can be a fundamental factor that determines the trust in the ML model. Deep learning models are complex and known to share “blackbox” nature, which raises a lot of concerns about their deployment in real-world applications despite their better performance compared to interestically-intereparable models such as decision trees and linear models. As a result, researchers have developed post-hoc explainers that identify a complex model’s behavior by analyzing its input, intermediate result, and output. A representative category in this vein approximates the non-linear decision surface either globally or locally by using an explainable ML model, i.e., an explainer, such as a linear model carvalho2019machine and rules guidotti2018local . For deep convolutional neural networks or transformers, the inspection of intermediate features is a widely used means of explaining model behavior tu2020select ; yang2018hotpotqa . In terms of robustness, our aim is sustainable predictive performance under data changes, especially, unexpected distribution shifts. However, predictive performance can include more fine-grained measures on the model’s behaviors than the accuracy of predictions. For instance, the post-hoc explanations obtained for novel inputs should also be stable, otherwise, the model is likely to become brittle due to changes in the data distribution. In the following, we will describe model testing coverage criteria that rely on intermediate states of a neural network to characterize its “behavior” when given a particular input. They can be seen as ways to capture the distribution shifts from the lens of the model itself by estimating how its intermediary states differ from the “normal” levels observed on the training samples.

### v3

@@ 1 Definition
In general, robustness is a predicate that applies to a single entity. For instance, we might consider a sensor robust if it is resilient to disturbances from the environment. In more detail, robustness refers to the ability of a system, model, or entity to maintain stable and reliable performance across a broad spectrum of conditions, variations, or challenges, demonstrating resilience and adaptability in the face of uncertainties or unexpected changes. Hence, we define below the Machine Learning (ML) Model robustness based on the general definition of robustness in ML outlined by Freiesleben and Grote in freiesleben2023beyond .
ML Model robustness denotes the capacity of a model to sustain stable predictive performance in the face of variations and changes in the input data.
According to this definition, ML Model robustness is an epistemic concept that presupposes the generalizability of the model’s inductive bias on the in-distribution data, and extends further to evaluate the model’s stability and resilience in real-world deployment scenarios. The following Table presents concrete illustrations of how performance degradation and data changes manifest in real-world scenarios.
[…]
Nevertheless, leaving the range of input data changes unspecified makes it hard to assess the robustness of ML model in practice. We should define the data changes against which the model would be tested. Even if the naturally occurring data distribution shifts are often unanticipated, we usually come up with data distortions that can serve as a proxy for unforeseen data shifts and help us compare the robustness of different ML models. Furthermore, the objective of sustaining stable predictive performance is vague. It is often sufficient for the model to maintain its performance to a certain degree (i.e., the tolerance level) against unexpected changes in input data. This level of tolerance depends on the application context and the assurances needed. For example, the target tolerance for an ML model designed to support clinical decision-making is considerably lower than for an ML model that has been designed to detect spam emails freiesleben2023beyond .
@@
1 Robustness complements (i.i.d.) generalizability
In supervised learning, models are commonly estimated via empirical risk minimization (ERM) goodfellow2016deep , a principle that considers minimizing the average loss on observed samples of data, as an empirical estimate of the true risk, i.e., the expected true loss for the entire input distribution. ERM assumes that training and test data are identically and independently distributed, known as closed-world assumption or simply the i.i.d. assumption. The i.i.d generalization refers to the ability of a trained model to deal with novel data inputs, but drawn from the same or close distribution as the training set, called in-distribution (ID) data. The i.i.d. generalization ensures stable predictive performance under static environmental conditions, but it provides no guidance on how to handle out-of-distribution data (OOD) yang2021generalized . In contrast, robustness focuses on capturing the level of predictive performance maintained by the trained model in dynamic environment settings, where input data constantly changes. Robustness could be of little concern if the model fails to i.i.d. generalize well. In order to achieve robustness, we consider i.i.d. generalization to be a necessary but not sufficient condition. For instance, ML models might fail to i.i.d. generalize due to unreliable inductive bias (shortcut learning) or under-fitting (trained with too little/biased data). These models will most likely also perform poorly when input data distributions change.
The ascent of DL has initiated a new era in AI, empowering models to tackle open-world learning challenges in domains like face recognition guo2016ms and autonomous driving bojarski2016end . The term “generalization” has expanded beyond denoting the model’s performance strictly in an i.i.d. setting, encompassing its ability to cope with OOD situations, hereafter referred to as OOD generalization. Although OOD generalization seems to fit the definition of robustness 1, it actually refers to the overall predictive performance of an ML model beyond its in-data distribution. It lacks a proper definition of its scope and success criteria since there is no indication of which data distributions the model should generalize to, or how strict the original predictive performance should be maintained. In contrast, ML model robustness is an inherently causal concept since it concerns two causally related entities: the level of predictive performance and the domain of input data changes, so its assessment, as a model’s property, requires detailed specification of both entities. Freiesleben and Grote freiesleben2023beyond provide further insight into the complementarity between robustness and generalizability.
@@ 2 Robustness is a requirement of Trustworthy AI
The adoption of ML, especially deep neural networks, has been largely promoted as a result of its impressive performance in terms of accuracy. Meanwhile, a variety of challenges outside of accuracy expectations have emerged, such as malicious attacks against ML-powered systems and misuses of ML that could be harmful. As a result, the Artificial Intelligence (AI) trustworthiness standards liu2022trustworthy ; li2023trustworthy have been established to outline representative requirements for current AI systems. These encompass six critical facets: (i) Safety & Robustness, (ii) Nondiscrimination & Fairness, (iii) Transparency & Explainability, (iv) Privacy, (v) Accountability & Auditability, and (vi) Environmental Well-being. The complex interplay between these aspects is vital in fostering trustworthy real-world AI systems. For instance, maintaining data privacy might interfere with the desire to explain the system output in detail.
@@
In this sense, robustness is an integral part of AI trustworthiness, while interacting and combining with the other aspects. In fact, model robustness is a cornerstone of safety because robust AI systems are able to deal with unexpected inputs and perturbations without compromising their functionality. It is especially crucial to ensure the appropriate level of model robustness in safety-critical applications, where erroneous behaviors and failures can have catastrophic consequences. The trustworthiness of AI systems requires the integration of two pivotal elements along with robustness assurance: reliable quantification of uncertainty and effective out-of-distribution detection capabilities.
Uncertainty quantification abdar2021review refers to methodologies for evaluating uncertainties associated with predictions made by a ML model. It involves assessing the confidence levels or lack thereof in the model’s predictions, taking into account factors such as data variance and model error. The quantified uncertainties combined with predictions enable more informed decision-making with AI.
Aleatoric uncertainty stems from inherent randomness in the data, such as noise, and is considered non-reducible. It reflects the unpredictability associated with the inherent variability within observations.
Epistemic uncertainty arises from limitations within the model’s bias and the training data used. It is considered reducible, as it can be mitigated through improvements in modeling strategy (e.g. ensembles), training procedures, or new data collection.
@@
In simpler terms, the uncertainty component provides the AI systems with a way to “know what they do not know”. Its main contribution is that uncertain predictions can be ignored from the decision-making flow, avoiding risks in real-world applications. Estimating uncertainties significantly benefits the model robustness analysis by providing an essential means to gauge how individual predictions respond to changes in their data points. This aids in delimiting the domain of input changes on which an ML model is expected to be robust.
@@
OOD detection yang2021generalized involves identifying OOD instances at test time that differ significantly from the in-data distribution and might result in mispredictions. It also serves the same purpose of recognizing the boundaries within which the model’s patterns are applicable, avoiding the use of its predictions when such restrictions are violated. OOD detection strengthens the reliability of uncertainty quantification by filtering out unusual inputs, on which the uncertainties, as any statistically-inferred estimates, are unlikely to be dependable. Therefore, OOD detection also contributes to further refinement of the data changes domain during the model robustness analysis, and its deployment in safety-critical systems ensures effective operation of robust ML models.
@@ 3 Adversarial vs Non-adversarial robustness
Non-adversarial robustness studies the model’s ability to maintain its performance across data distribution shifts arising from naturally-occurring distortions or synthetic data variations that represent conditions more likely to occur in the real world. For instance, a natural shift in images of traffic signs, collected in an area where it rarely snows, can be images of the same sign under severe snowing conditions gojic2023non . Alternatively, a partial discoloration of the traffic sign image, i.e., a region replaced by white pixels, may mimic the effect of snow or other neutral obstructions. Natural data shifts often result from changing environmental conditions that lead to mismatch between the deployment and the training distribution xie2019multi . The changes can be both temporal (i.e., changes in patient demographics, genetic evolution of contagious viruses and their adaptation to host populations) or non-temporal (i.e., variations between patient groups, differences in imaging technologies used by diagnostic devices). The non-adversarial robustness of the ML model ensures reliability across various real-world scenarios, including natural noise, or changes in inputs that might happen organically without malicious intent.
@@ 4 Robustness Measurement
We now present several robustness scoring metrics developed by researchers to evaluate the stability of models’ performance when data changes drenkow2021systematic . These metrics are designed with the assumption of having two datasets: one clean, sampled from ID data, and another perturbed, sampled from shifted or even OOD data. Specifically, shifted data involves significant variations and unusual scenarios within the same domain as the ID data, while OOD data represents inputs from domains that are fundamentally different from the ID data.
@@
The more the robustness score of a model is close to one, the more it is robust to the considered perturbation. To measure the robustness score of a neural network, $f$, to a set of perturbations $S$, we can use:
$R_{f}^{S}=\sum_{\phi\in S}R_{N}^{\phi}$
Hendrycks and Dietterich hendrycks2019benchmarking define robustness as average-case performance over a set of corruptions, leading to the definition of mean Corruption Error (mCE) and relative Corruption Error (rCE).
@@
where $E$ is the classifier’s error rate, $\mathcal{C}$ is a set of corruptions, and $s$ is the severity of the corruption (refer to Section 1 for examples). The AlexNet model krizhevsky2012imagenet served as a common reference point among models.
Effective and relative robustness taori2020measuring consider how the performance on natural distribution shifts related to the performance on the original test set. The metrics are defined as follows:
$aligned\rho(f)&=acc_{2}(f)-\beta\cdot acc_{1}(f),\\ \tau\left(f^{\prime}\right)&=acc_{2}\left(f^{\prime}\right)-acc_{2}(f)$
where $f$ is the model under test, $acc_{1}$, $acc_{2}$ are the accuracy on the original and shifted datasets respectively, $\beta$ is a log-linear fit to the baseline accuracy of a large set of independent models on the original (clean) test set, and $f^{\prime}$ represents the model resulting from a robustness improvement. The effective robustness captures how well a specific model does beyond what is expected given a group of models in general. The relative robustness quantifies the effect of a robustness amelioration strategy on the accuracy under a distribution shift. Overall, a robust model should obtain both positive effective and relative robustness scores.
@@ 5 Robustness Indicators: Reproducibility and Explainability
So far, robustness has been considered as a quality predicate for an ML model that should be carefully specified, evaluated, and sometimes certified before the deployment of the model in production. Nevertheless, the ML model is the result of a ML engineering process, thus some desirable properties in regards to this process should be satisfied in order to prepare the grounds to reach the desired level of robustness.
@@
Reproducibility controls the randomness of an ML workflow, making it resilient to slight variations in data samples, i.e, there are high chances to converge to the same conclusions when run on two samples from the same underlying distribution albertoni2023reproducibility . For researchers in all scientific fields, including ML, reproducibility is essential to achieve the same results with the same data and algorithms. The absence of reproducibility may lead to claim gains from changing one parameter while the real source of improvement is a hidden source of randomness. Reproducibility is also a strong stability property since it ensures a high probability of replicating results when the datasets are drawn from the same distribution. Consequently, achieving such property on a given learning problem ensures less bias in the estimated model, $f$, since achieving such property confirms that the difference between the empirical risk of $f$ obtained on the training samples and the (true) risk of $f$ is marginal, which implies that M is more likely independent of the training set. For instance, underspecificationMLmodel constructs an ensemble of predictors from a given model by perturbing components of the ML pipeline like random seeds for initialization, and retraining the model several times. As a result, output differences between predictors against stress tests can be used as a conservative indicator of the ML pipeline credibility, i.e., its ability to produce a robust model.
@@
Explainability addresses how an AI model makes decisions carvalho2019machine . Being aware of the reasoning behind predictions can be a fundamental factor that determines the trust in an ML model. DL models are complex and known to share “blackbox” nature, which raises concerns about their deployment in real-world applications despite their better performance compared to intrinsically-interpretable models such as decision trees and linear models. Researchers have developed post-hoc explainers that identify a complex model’s behavior by analyzing its inputs, intermediate results, and outputs. In complex models, a nonlinear regression function or nonlinear classification boundary is learned. A representative category of post-host explainers approximates these non-linear decision functions either globally or locally by using an explainable ML model, i.e., an explainer, such as a linear model carvalho2019machine and rules guidotti2018local . For deep convolutional neural networks or transformers, the inspection of intermediate features is a widely used means of explaining model behavior tu2020select ; yang2018hotpotqa . As an example from image classification, saliency maps simonyan2013deep indicate how much influence each input feature (i.e. pixel) has on predicting a specific class. With color coding for their value ranges, they can visually show the regions of interest in the image that have the greatest impact on the classification result. Arun et al. arun2021assessing leverages a variety of saliency map techniques to interpret DNNs trained on medical imaging. In terms of robustness, our aim is sustainable predictive performance under data changes, especially, unexpected distribution shifts. Post-hoc explainers can provide other predictive performance indicators on the model’s behaviors in addition to the accuracy of predictions. Indeed, inconsistent post-hoc explanations for novel inputs when compared with ID samples may indicate erroneous model behaviors due to changes in the data distribution. For instance, saliency maps can generate adversarial noise JSM affecting the most significant regions of interest, and ML testing methods braiek2020testing can be used to find edge cases that lead the model to mistakenly rely on irrelevant regions. They can be seen as ways to capture the distribution shifts from the lens of the model itself by estimating how its behavior (e.g., as captured by post-hoc explanations) differ from the “normal” behavior observed on the ID samples.